OpenStudy (anonymous):

How to simplify 1^16 sqrt(x*x^3*x^4)?

OpenStudy (anonymous):

\[1^{16} \sqrt{(x*x^3*x^4)}\] is this what you mean???

OpenStudy (anonymous):

Yes, except the 1 isn' there, just the ^16. I dunno if that makes a difference or not. :) @azetina

OpenStudy (anonymous):

So is it \[\sqrt[16]{x*x^2*x^4}\]

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

Only x^3 instead of x^2 :)

OpenStudy (anonymous):

@azetina

OpenStudy (anonymous):

\[\sqrt{x}\] is the answer

OpenStudy (anonymous):

Can you show your work? wolframalpha gives me sqrt(x^8)

OpenStudy (anonymous):

I need to go and teach a math of finance class be back later if you dont mind

OpenStudy (anonymous):

..okay..

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg <-- notice the 1st rule listed there

OpenStudy (jdoe0001):

\(\large \bf 1^{16}\sqrt{x\cdot x^3\cdot x^4}\qquad \qquad 1^{16}=1\qquad x^a\cdot x^b\cdot x^c=x^{a+b+c}\)

OpenStudy (jdoe0001):

same base, different exponents, when multiplying, keep the base, add the exponents

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw